Two bodies of masses 2kg and 4kg are moving with velocities 2m/s and 10m/s towards each other due to mutual gravitational attraction.then the velocity of the centre of mass is
 

Grade:11

3 Answers

Khimraj
5 years ago
velocity of COM = (m1v1 + m2v2)/(m1+m2) = (40-4)/6 = 6 m/s
Hope it clears...............................................................................
Simran
5 years ago
Thanks sir, but in a book its answer is zero and the reason behind it is that centre of mass remains at rest under mutual gravitational attraction
Rahul
4 years ago
2kg mass ,  and 4 kg mass. 
V1=20m/s and v2=10m/s 
 Now,
               Vcom=2*20+4(-10)/6
       then , Vcom = 40-40/6
                             =0
So the answer zero is correct